# Improving Human-AI Coordination through Online Adversarial Training and Generative Models

### v3

@@ 7 Conclusion
Our work demonstrates that combining generative modeling with adversarial training provides an effective approach for training AI agents that can coordinate with diverse human partners. Using a generative model and regret-based optimization to constrain the Adversary enables GOAT to generate valid, meaningful cooperation partners while posing a curriculum of increasingly challenging scenarios, adapted to the weaknesses of the Cooperator’s policy. Thus, we enable robust learning without the risk of degenerate strategies. Experimental results on the Overcooked benchmark show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners. The improved coverage of the partner space demonstrated by our agents suggests this approach successfully bridges the gap between training efficiency and the generalization to human behavior. Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training.
8 Acknowledgements
This work was generously supported by the Cooperative AI Foundation, UW-Amazon Science Gift Hub, Amazon Middle Mile Products and Technology Team, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), and Microsoft Accelerate Foundation Models Research Program.
@@ Appendix A Reproducibility
We have our live interactive demo and extended experimental data on this webpage. https://sites.google.com/view/goat-2025. We provide all the details to set up the environment and reproduce the results. For the One-Step Cooperative Matrix Game please find Google Colab with code and all the instructions to run the code.
@@ Appendix B Implementation details
Cooperator Agent The Cooperator agent was trained using PPO [29]. To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up. We adapted our codebase on the previous implementation of HSP [41] and GAMMA [18].
GOAT For the GOAT policy training, we use the REINFORCE objective with KL divergence penalty for drifting away from the prior, mean, and standard deviation hard bounds to keep the policy constrained in the VAE latent space. This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
@@ Appendix C Hyperparameters
The GOAT uses a simple policy and was trained using REINFORCE to generate latent vectors. The architecture and hyperparameters are defined below.
@@
output dimension 2×2$\times$2 × z_dim
@@
KL coefficient 0.1
entropy coefficient 0.01
std scaling 4.0, 1.0
mean scaling 60.0
gradient clipping 1.0
learning rate scheduler start factor 1.0
learning rate scheduler end factor 0.9
@@ Appendix D Computational Resources
Our primary tests were carried out on AMD EPYC 64-Core Processor and NVIDIA L40s/L40 clusters. One Cooperator agent can be trained in roughly a day. The primary tests require roughly 3600 GPU hours. We perform some preliminary experiments to determine the best training frameworks and hyperparameters.
Appendix E Minimax and Regret comparison on overcooked
Figure 7: Performance comparison between minimax vs GOAT objective on Multi-Strategy Counter when evaluated against held-out human behavior cloned model as partner. Error bars are the Standard Error of the Mean.
Appendix F Cooperative strategies learned by VAE
Appendix G Bonferroni Tests
Hypothesis Counter circuit ($p$-value) Multi-Strategy Counter ($p$-value)
CoMeDi $1.93\times 10^{-11}$ $9.50\times 10^{-181}$
FCP $1.71\times 10^{-47}$ $2.24\times 10^{-53}$
BC $+$ RL $5.19\times 10^{-15}$ $1.20\times 10^{-20}$
MEP $1.0$ $3.48\times 10^{-28}$
MEP $+$ GAMMA $0.349$ $1.31\times 10^{-16}$
[…]
Appendix H Tests on CMG-S
@@ Appendix I Broader Impact Statement
Our work on human-AI coordination aims to provide humanity with better cooperative AI that can adapt, understand, and generalize to diverse humans. While this could enhance collaborative robots, autonomous vehicles, and assistive technologies, our method specifically improves generalization to new partners without requiring extensive human data. GOAT has the potential to improve capabilities for household robots, vehicles, and collaborative AI systems where seamless human interaction adds value to people’s lives. This type of technology enhances human modelling, which can be deployed widely to help people from all walks of life and of all age groups. We would also like to note that the development of such an AI system will accelerate technological progress and requires responsible development and deployment.
Appendix J Human qualitative ratings.
(a) Counter circuit layout
(b) Multi-strategy Counter
Figure 11: Human evaluations of agent behavior. Participants rated their agreement with the statement "The agent adapted to me when making decisions" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree. GOAT shows decent adaptability on simpler layout, whereas it appears more adaptable on harder layout.
(a) Counter circuit layout
(b) Multi-strategy Counter
Figure 12: User assessment of agent coordination. Participants rated their agreement with the statement "The agent’s actions were human-like" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree. GOAT is seen as more confident in coordinating on harder layouts.
(a) Counter circuit layout
(b) Multi-strategy Counter
Figure 13: User assessments of agent frustration levels. Participants rated their agreement with the statement "The agent’s behavior was frustrating" using a 5-point Likert scale ranging from Strongly Disagree to Strongly Agree. GOAT receives comparable ratings for showing less frustrating behavior compared to their respective baselines.
Appendix K Limitation
The effectiveness of this method heavily depends on the training of VAE models and enough data available for VAE to learn the distribution instead of overfitting to a few examples. It would be a promising method for fields like robotics and human-AI interaction, where data availability is not an issue. We also found that in sparse reward problems, this method discovers solutions and trains the cooperator robustly.

### v4

@@ 7 Conclusion
Our work demonstrates that combining generative modeling with adversarial training provides an effective approach for training AI agents that can coordinate with diverse human partners. Using a generative model and regret-based optimization to constrain the adversary enables GOAT to generate valid, meaningful cooperation partners while posing a curriculum of increasingly challenging scenarios, adapted to the weaknesses of the Cooperator’s policy. Thus, we enable robust learning without the risk of degenerate strategies. Experimental results on three cooperation tasks show that GOAT achieves state-of-the-art performance in both simulated evaluations and studies with real human partners. The improved coverage of the partner space demonstrated by our agents suggests this approach can successfully and robustly achieve generalization to human behavior. Future work could explore extending this framework to more complex collaborative tasks and investigating ways to incorporate explicit human feedback during training.
8 Acknowledgments
This research was supported by the Cooperative AI Foundation, the UW-Amazon Science Gift Hub, Sony Research Award, UW-Tsukuba Amazon NVIDIA Cross Pacific AI Initiative (XPAI), the Microsoft Accelerate Foundation Models Research Program, Character.AI, DoorDash, and the Schmidt AI2050 Fellows program. This material is based upon work supported by the Defense Advanced Research Projects Agency and the Air Force Research Laboratory, contract number(s): FA8650-23-C-7316. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of AFRL or DARPA.
@@ Appendix A Reproducibility
We have our live interactive demo, extended experimental data, and code on our webpage, link on first page.
@@ Appendix B Implementation details
Cooperator Agent The Cooperator agent was trained using PPO (Schulman et al., 2017). To promote exploration, the first 100M steps use reward shaping for dish and soup pick-up. We adapted our codebase on the previous implementation of HSP (Yu et al., 2022) and GAMMA (Liang et al., 2024).
GOAT For GOAT policy training, we use the REINFORCE objective with KL divergence penalty to keep the policy constrained in the VAE latent space. This objective requires two separate passes of the environment to gather returns for self-play and cross-play.
@@ Appendix C Hyperparameters
GOAT uses a simple policy and was trained using REINFORCE to generate latent vectors. The architecture and hyperparameters are defined below.
@@
output dimension $2\times$ z_dim
@@
KL coefficient 5.0
@@ Appendix D Computational Resources
Our primary tests were carried out on AMD EPYC 64-Core Processor and NVIDIA L40s/L40 clusters. One Cooperator agent can be trained in roughly a day. The primary tests require roughly 24-36 GPU hours for Overcooked and 1 hour for rest of the environments. We perform some preliminary experiments to determine the best training frameworks and hyperparameters.
Appendix E Cooperative Reaching Game: Game layout & 11 Heuristic Agents
Figure 6: Cooperative Reaching Game (CRG): Red and Blue are the two agents that should cooperate to reach and stay on one of the goal coordinates in the corner. Yellow goal coordinates lead to a reward of 11, and Green goal coordinates lead to a reward of $0.75$.
  • Heuristic H01: Selects actions moving teammates toward the nearest reward-providing coordinate.
  • Heuristic H02: Selects actions moving teammates toward the reward-providing coordinate most distant from their episode starting position.
  • Heuristic H03: Moves teammates toward the nearest optimal reward-providing coordinate.
  • Heuristic H04: Moves agents toward the optimal reward-providing coordinate most distant from teammates’ initial episode location.
  • Heuristic H05: Same as H4, but considers only suboptimal reward-providing coordinates instead of optimal ones.
  • Heuristic H06: Same as H5, but teammates move toward the nearest suboptimal reward-providing coordinate.
  • Heuristic H07: At episode start, agents randomly select a reward-providing coordinate and move toward it.
  • Heuristic H08: Moves teammates toward the reward-providing coordinate nearest to their counterpart agent’s location.
  • Heuristic H09: Same as H8, but considers only optimal reward-providing coordinates as destinations.
  • Heuristic H10: Moves teammates directly toward their counterpart agent’s location.
  • Heuristic H11: Always randomly selects actions from teammates’ available action set.
Appendix F Tests on CMG-S
@@
Appendix G Minimax and Regret comparison on overcooked
Figure 8: Performance comparison between minimax vs GOAT objective on Multi-Strategy Counter when evaluated against a held-out human behavior cloned model as a partner. Error bars are the Standard Error of the Mean. GOAT outperforms MiniMax, but here MiniMax produces poor-performing agents that fail to participate in coordination. See game videos on our website.
Appendix H Cooperative strategies learned by VAE
@@ Appendix I Broader Impact Statement
Our work on human-AI coordination aims to provide humanity with better cooperative AI that can adapt, understand, and generalize to diverse humans. While this could enhance collaborative robots, autonomous vehicles, and assistive technologies, our method specifically improves generalization to new partners without requiring extensive human data. GOAT has the potential to improve capabilities for household robots, vehicles, and collaborative AI systems where seamless human interaction adds value to people’s lives. This type of technology enhances human modeling, which can be deployed widely to help people from all walks of life and of all age groups. We would also like to note that the development of such an AI system will accelerate technological progress and requires responsible development and deployment.
Appendix J Limitation
The effectiveness of this method depends on the training of VAE models and enough data available for VAE to learn the distribution, instead of overfitting to a few examples. It would be a promising method for fields like robotics and human-AI interaction, where data availability is not an issue. We also found that in sparse reward problems, this method discovers solutions and trains the cooperator robustly.